OpenStudy (anonymous):

how much will the 5% on 6000 euros cost me

OpenStudy (anonymous):

5% of 6000 means: \[\frac{5}{100} \times 6000\] Can you solve it now??

mathslover (mathslover):

x% means x/100 so we have : 5 % = 5/100 just think that if u have 100 toffees how many will left after u ate 5 of them..

mathslover (mathslover):

\[\large{5 /100 *6000 = \frac{5}{100}*6000 = \frac{30 000}{100} = 300}\]

mathslover (mathslover):

like this..

OpenStudy (anonymous):

@emceejay you are expected to interact here..

mathslover (mathslover):

right... please @emceejay ...

OpenStudy (anonymous):

Anyways I have a surprise for you that @mathslover taught me..

mathslover (mathslover):

:P what did i teach you @waterineyes ?

OpenStudy (anonymous):

Wait I show you what you have taught me..

OpenStudy (qqstory):

shouldn't 0.05 times 6000 be easier than what u guys said..?

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy...}}\]

mathslover (mathslover):

haha

OpenStudy (anonymous):

Yeah you can look at this that way also @emceejay ..

OpenStudy (anonymous):

But for calculation I think @mathslover is right..

OpenStudy (anonymous):

\[0.05 \times 6000 = \frac{5}{100} \times 6000 = 5 \times 60 = 300\]

mathslover (mathslover):

thanks (where is br then )? : P

OpenStudy (qqstory):

or easier? u just move the 0.05 to 5 and remove 2 zero from 6000, which means u have 60 times 5.. = 300?

OpenStudy (qqstory):

well i am better with decimals in this case, so its really up to the person LOLOLOL.

OpenStudy (lgbasallote):

if three people are talking at the same time..who are you supposed to respond to? hmm